OpenStudy (anonymous):

find the indicated matrix 3E [0 2 ] 1 -4

OpenStudy (anonymous):

I dont get it....can you give us more information

OpenStudy (anonymous):

If the question is find 3E where E=\(\begin{bmatrix}0&2\\1&-4\end{bmatrix}\), then it's just scalar multiplication of a matrix, which is simply 3E=\(\begin{bmatrix}3\cdot0&3\cdot2\\3\cdot1&3\cdot-4\end{bmatrix}\). I'm not sure if that's what you're asking or not though.

OpenStudy (anonymous):

im not either but ill just look it over again

OpenStudy (anonymous):

what @nbouscal said. multiply every entry by 3

OpenStudy (anonymous):

oh okay thanks